OpenStudy (anonymous):

12=21-3(11) gets a negative 12 not 12? How do I change that!?

OpenStudy (anonymous):

21 - 3(11) = 21 - 33 = 21 - (21 + 12). You should expect to get a negative answer since you're subtracting more than 21 away from 21. 21 - (21 + 12) = 21 - 21 - 12 = - 12. You should get -12.